OpenStudy (anonymous):

If a number with 3 digits divided by 13 gives a certain number(y) +11, divided by 11 gives a certain number (z)+ 9 and divided by 9 gives a number (q) +7, what is the value of the 3 digit number? My English might not be perfect and stuff but gosh, halp please been trying for the last 2 hours

OpenStudy (anonymous):

a number with 3 digits divided by 13 gives a result +11 should be as simple as 13 * 11 = 143

OpenStudy (anonymous):

9*11 = 99 9*7 = 63

OpenStudy (anonymous):

but i think the problem is asking for the same # for all 3 problems...

OpenStudy (anonymous):

Yeah, the x Is supposed to be the same for all 3 of them.. which leads me to a problem with 4 variables that I can't solve for the life of me

OpenStudy (anonymous):

I really do not see 4 variables..what topic of math is this for?

OpenStudy (anonymous):

I'm not sure to be honest, I think it was some way for our teacher to see how far we've gotten with our studies. I might not have been clear though, by "A result +11" I meant the result as in a number + 11, not that the result is 11. Sorry, I'll edit that now What I got with 4 variables was: X/13 = Y + 11 X/11 = Z + 9 x/9 = Q + 7

OpenStudy (anonymous):

Actually, X=13A+11=11B+9=9C+7 where A,B and C are integers

OpenStudy (anonymous):

Since 9,11 and 13 has no common factors this question should be too hard. Let X = 13*11*9*A + 13*11*B + 13*9*C + 11*9*D

OpenStudy (anonymous):

Now, we divide X by 13 we should get 11 as remainder So, X/13= 11*9A +11*B +9*C + (11*9*D/13)

OpenStudy (anonymous):

(99 *D )mod 13=11 we need to solve this to find D

OpenStudy (anonymous):

Similarly , (13*11*B) MOD 9 =7 (13*9*C) MOD 11=9

OpenStudy (anonymous):

So, you should get after solving that B=2 , C=6 and D=3

OpenStudy (anonymous):

So, X=13*11*9*A + 13*11*B + 13*9*C + 11*9*D =1287*A + 143*2+117*6+99*3 =1287*A+1285

OpenStudy (anonymous):

But I am not getting any three digit number.... 1285 looks like is the smallest positive integer

OpenStudy (anonymous):

Thank you! The way you solved it was a lot different than the way our teacher did though, she solved it with a gimmick 8D And she was mistaken about the 3 digit number, the correct number was 1285.